OpenStudy (anonymous):

De Moivre's theorem: Use de Moivre's Theorem to find the following expression. Write your answer in standard form. (-sqrt3 + i)^4... someone look over my work and tell me what I am doing wrong??!!

OpenStudy (anonymous):

The answer is pizza, it always has been it always will be

OpenStudy (anonymous):

= 2(cos 30 + i sin30)^4 = 16(cos120 + i sin120) = 16(-1/2 + sqrt3 /2 i) = -8+8sqrt3i

OpenStudy (anonymous):

IT WILL ALWAYS BE PIZZA

OpenStudy (anonymous):

ALWAYS

OpenStudy (anonymous):

jim_thompson5910 can u look over this? please?

jimthompson5910 (jim_thompson5910):

notice how we have -sqrt(3) and 1 from -sqrt3 + i = -sqrt3 + 1i if you divide each by 2 to get -sqrt(3)/2 and 1/2. Those two values are found on the unit circle

jimthompson5910 (jim_thompson5910):

Let's make z = -sqrt(3) + i divide everything by 2 to get z/2 = -sqrt(3)/2 + (1/2)i

jimthompson5910 (jim_thompson5910):

then compare -sqrt(3)/2 + (1/2)i to cos(theta) + i*sin(theta) what must the value of theta be?

OpenStudy (anonymous):

wow hold on i lost u

jimthompson5910 (jim_thompson5910):

sorry about that where are you stuck?

OpenStudy (anonymous):

ok why would you divide it by two?

jimthompson5910 (jim_thompson5910):

-sqrt(3) is not found on the unit circle but -sqrt(3)/2 is found on the unit circle

ganeshie8 (ganeshie8):

dividing by 2 and comparing lets you avoid calculator in computing the angle

OpenStudy (anonymous):

ok, can someone write it out for me? because I am not getting it... You cant just randomly divide two to avoid computing, so I am sure you are saying something that makes a lot of sense to you guys but for me is in another language

ganeshie8 (ganeshie8):

are you allowed to use calculator ?

ganeshie8 (ganeshie8):

if so find the angle of \(a+ ib \) by the usual formula \[\theta = \tan^{-1}\left(\frac{b}{a}\right)\]

ganeshie8 (ganeshie8):

the angle you have calculated in the start was wrong, work it again

OpenStudy (anonymous):

i did ... isnt it 120?

OpenStudy (anonymous):

ok let me see...

ganeshie8 (ganeshie8):